A girl have 5 sister having sum of there age is 90, and diffrence of there age is 3. Find out the age of eldest sister.?
Answer by ankor@dixie-net.com(22740)

A girl have 5 sisters having sum of there age is 90, and difference of there age is 3. Find out the age of eldest sister.?
:
Assume there is one sister plus one set of twins and one set of triplets
g = girl's age, who is the oldest
(g-2) = the triplets age
(g-3) = the twin's age
the difference between the oldest and the youngest will be 3
:
g + 3(g-2) + 2(g-3) = 90
g + 3g - 6 + 2g - 6 = 90
6g - 12 = 90
6g = 90 + 12
g = 102/6
g = 17 yrs is the age of the oldest
:
:
See if that works, the triplets will 15, the twins will be 14
17 + 3(15) + 2(14) =
17 + 45 + 28 = 90
:
I have to admit I had to try several combinations of twins and triplets to get an integer solution to this